Session: Feb 2000 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Juri Roda at Tue Feb 29 04:00:44 MET 2000) Success
No problem

Onsala (A. Polatidis at Tue Feb 29 08:41:39 MET 2000) Minor failures
Observed until 2023 UT. The antenna was then stowed due to strong winds. We are sending the first 3 tapes only.

Torun (K.Borkowski at Tue Feb 29 09:22:28 MET 2000) Success
Strong wind during the experiment may have affected the pointing. No other known problems.

Westerbork (Tony Foley at Tue Feb 29 14:48:36 MET 2000) Minor failures
Polarizations swapped from 11:05 to 11:11UT Late tape change at 14:45 lost scans from 14:35 to 14:58 Tape broke at 21:25 - had to restart schedule at 21:35. From 17:30 water dripping into console room from new roof!

Effelsberg (A. Kraus at Wed Mar 1 09:47:39 MET 2000) Minor failures
Two scans lost (UT1138-1145) - problems with the telescope control computer.

Jodrell (Alastair Gunn at Thu Mar 2 05:35:45 MET 2000) Minor failures
Telescope parked at about 11:45 UT due to high winds, back on source at about 13:07 UT. No other known problems.

Cambridge (Alastair Gunn at Thu Mar 2 05:36:17 MET 2000) Success
No known problems.

Noto (Noto VLBI friend at Thu Mar 2 12:07:59 MET 2000) Success

Calibration (Alastair Gunn at Wed Apr 19 13:09:19 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/er008b.antab for details